You would like to buy a house, but you can only afford monthly payments of R500. Thirty-year loans are available at 6% interest compounded monthly. If you can make a R5 000 down payment, what is the price of the most expensive house that you can afford to purchase?


1
Expert's answer
2022-04-01T11:50:58-0400

let's find an annuity


"\frac{0.06}{12}=0.005"


"FV=500(\frac{1+0.005)^{360}-1}{0.005})=502257.52"


the price of the most expensive house that you can afford to purchase:

502257.82+5000=507257.52